OpenStudy (sissyedgar):

In an election, 11 of 20 voters preferred Ms. Williamson to Ms. Stallings. A total of 3300 people voted in the election. How many voters preferred Ms. Williamson to Ms. Stallings? A. 1815 B. 3291 C. 4185 D. 6000

OpenStudy (kohai):

So if it's a statistic out of 20, and 3000 participated, you have to divide 3000 by 20 to get what you have to multiply by. Then you multiply that number by the 11

OpenStudy (sissyedgar):

lol, i think u ment 3300

OpenStudy (kohai):

 If that makes sense Yeah I misread lol

OpenStudy (sissyedgar):

A.

OpenStudy (kohai):

Yeah, good job :)

OpenStudy (kohai):

And if you want to check your work, divide both sides by 165 and you'll get 11/20

OpenStudy (sissyedgar):

can u help wit another?

OpenStudy (sissyedgar):

@kohai @texaschic101

OpenStudy (sissyedgar):

I think D!

OpenStudy (sissyedgar):

@misty1212

OpenStudy (misty1212):

\[3(x+2y)=5x\] is a start

OpenStudy (misty1212):

so \[3x+6y=5y\\ 3x=-y\]

OpenStudy (sissyedgar):

okay so its either A. or B. rite?! gurl i been missin u! i neede help earlier. it was hard 2 trust sum1 2 get meh 2 pass.

OpenStudy (misty1212):

lets finish it

OpenStudy (misty1212):

\[\frac{-3x}{y}=1\\ \frac{x}{y}=-\frac{1}{3}\]

OpenStudy (sissyedgar):

o, ok.. thx @misty1212